# Quasiconvexity in the Riemannian setting

### v1

@@ 1. Introduction
is sequentially weakly lower semicontinuous in $W^{1,1}(a,b)$ if and only if the function $f$ is convex in the third variable. In the scalar case $m=1$, this result was later generalized to functions defined on bounded open sets of $\mathbb{R}^{n}$ by various authors and Serrin in [12] proved that the differentiability assumptions are not actually required. Improvements of Serrin’s theorem have then been given by De Giorgi [3], Olech [11] and Ioffe [6].
@@
In the case of $m>1$, the condition on $f$ that turned out to be necessary and, with additional assumptions, also sufficient for the lower semicontinuity of integral functionals, is the quasiconvexity, introduced by Morrey [10] in 1952.
@@
A real function $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{{n\times m}}\to\mathbb{R}$ is quasiconvex in $\xi\in\mathbb{R}^{n}$ if there exists a subset $Z$ of $\mathbb{R}^{n}$ with ${{\mathscr{L}}}^{n}(Z)=0$, such that for every ${x}\in\mathbb{R}^{n}\setminus Z$ and for every ${s}\in\mathbb{R}^{m}$ the function $\xi\mapsto f({x},{s},\xi)$ is quasiconvex.
More precisely, Morrey showed that under some strong regularity assumptions on the function $f$, the equivalence between its quasiconvexity and the weakly sequential lower semicontinuity in $W^{1,\infty}(\Omega,\mathbb{R}^{m})$ of the functional
@@ 2. Quasiconvexity in the Riemannian setting
In this section, after introducing a suitable “Riemannian” definition of quasiconvexity generalizing the “classical” Euclidean one, we will prove that the weak sequential lower semicontinuity in $W^{1,\infty}$ of a functional holds if and only if the integrand is quasiconvex, according to such definition.
@@
(where the identification of $T_{x}M$ with ${{\mathbb{R}}}^{n}$ is done via an orthornormal basis of $T_{x}M$) and its diffeomorphic image
@@
We chose to adopt the notation $du[x]$ for the differential of a map $u:M\to{{\mathbb{R}}}^{m}$ at a point $x\in M$, for sake of clarity in the computations that follow.
@@
Let $(M,g)$ be a smoooth, connected and complete Riemannian manifold and $\mu$ its canonical volume measure. A continuous function $f:\mathscr{L}(TM,\mathbb{R}^{m})\to\mathbb{R}$ is called quasiconvex if for every $x_{0}\in M$, $\alpha_{x_{0}}\in(T_{x_{0}}M)^{m}\subseteq\mathscr{L}(TM,\mathbb{R}^{m})$ and $\varphi\in C^{\infty}_{c}(Q_{x_{0}}^{r},\mathbb{R}^{m})$, there holds
@@
is sequential lower semicontinuous in the weak* topology of $W^{1,\infty}(\Omega,{{\mathbb{R}}}^{m})$, hence $\widetilde{f}$ is quasiconvex according to the Definition 1.1, by the Acerbi–Fusco Theorem 1.2.
@@ 3. Quasiconvexity and semicontinuity
We are going to show the analogues of some special cases of Theorem 1.2 of Acerbi and Fusco in our context, following the line of [2] (we mention that the main argument in the next Theorem 3.1 is due to Fonseca–Muller [5]). We underline (see the discussion above) that we are generalizing the special case of integrands which are continuous and depends only on $x$ and $Du$ (but not on $u$ – see the final section), from the Euclidean to the Riemannian setting.
@@
Let $n\in{{\mathbb{N}}}$ be the dimension of $M$. Let $x_{0}\in\Omega$ and we assume that $r>0$ is small enough such that $Q_{x_{0}}^{r}$ is contained in a neighbourhood $U\subseteq\Omega$ of $x_{0}$ such that the vector bundle $W^{1,\infty}(\Omega,{{\mathbb{R}}}^{m})$ can be “trivialized” as we discussed above at the beginning of this section, by choosing an orthonormal frame $E_{1},\dots,E_{n}$ in $U$, with an associated map $I$ and distance $\delta$ as in formulas (2.1) and (2.2). Moreover, since in what follows all the arguments of the continuous function $f$ will be bounded (since we are working in $W^{1,\infty}(\Omega,{{\mathbb{R}}}^{m})$ and $u_{j}$ is a bounded sequence in $W^{1,\infty}$, being weakly convergent), we can assume that $f$ is bounded and has a uniform modulus of continuity $\omega$ in $\mathscr{L}(TU,\mathbb{R}^{m})$ (with respect to the distance $\delta$) which is continuous, bounded and concave, hence subadditive. Moreover, since it is bounded, by possibly adding a constant to $f$, we can also assume that it is positive.
@@
where the function $o(1)$ is independent by $j\in{{\mathbb{N}}}$, while the functions $o_{j}(1)$ are equal to $\eta(r,\|d\varphi_{j}\|)$ for some function $\eta$ going to zero as $r\to 0$ and monotone nondecreasing (and we can also clearly assume continuous from the right) in its second argument.
@@
$\varphi_{h}\mathrel{\hbox{\set@color\hskip 5.83331pt\hskip-5.83331pt\hbox{\set@color$\xrightharpoonup{\ \ \,}$}\hskip-5.83331pt\hskip-3.71524pt\raisebox{-2.5pt}{\hbox{\set@color$\scriptstyle{\,\star\,}$}}\hskip-3.71524pt\hskip 5.83331pt}}0\quad\text{ in }W^{1,\infty}_{0}(Q_{x_{0}}^{r},\mathbb{R}^{m})$
@@
for every $r>0$ small enough. Then, arguing as in the proof of Theorem 3.1 (more precisely, following the same argument leading to estimate (3.13)), considering a suitable modulus of continuiy $\omega$ of $f$, we have
@@ 4. Some remarks and possible research directions
A natural continuation of our work would consist in extending the previous results also to the $W^{1,p}(\Omega,\mathbb{R}^{m})$ setting, with $p\in[1,+\infty)$, in order to obtain the analogue of Theorem 1.3 (notice that the fact that the semicontinuity of the functional implies the quasiconvexity of $f$ follows immediately by Theorem 3.2 in such setting, as in the Euclidean case).
Focardi and Spadaro in [4] showed an extension of the Acerbi–Fusco theorems (with $f$ continuous, like us) in the case of Sobolev maps $u:\Omega\to M$, with $\Omega$ an open bounded subset of $\mathbb{R}^{n}$ and $M$ a Riemannian manifold (this case actually “correspond” to the Euclidean situation with a continuous function $f$ depending on $u$ and $Du$, but not on $x$). Therefore, a possible future research line could be to “combine” Theorem 3.3 with the results of Focardi and Spadaro, thus obtaining a characterization of quasiconvexity in the completely Riemannian context of Sobolev maps between two Riemannian manifolds.
Moreover, the full generalization of the Acerbi–Fusco results would be to show them for the “naturally defined” Carathéodory functions on a Riemannian manifold, that is, functions which Carathédory in the usual way after expressing them by means of the local “trivializations” of the bundle $\mathscr{L}(TM,\mathbb{R}^{m})$ (or of $\mathscr{L}(TM,TN)$ after “combining” our definition with the one of Focardi–Spadaro), as we did at the beginning of this section.

### v2

@@ 1. Introduction
is sequentially weakly lower semicontinuous in $W^{1,1}(a,b)$ if and only if the function $f$ is convex in the third variable. In the scalar case $m=1$, this result was later generalized to functions defined on bounded open sets of $\mathbb{R}^{n}$ by various authors and Serrin in [12] proved that the differentiability assumptions are not actually required. Subsequent improvements of Serrin’s theorem were given by De Giorgi [3], Olech [11] and Ioffe [6].
@@
In the case of $m>1$, the condition on $f$ that turned out to be necessary and, with additional assumptions, also sufficient for the lower semicontinuity of integral functionals, is quasiconvexity, introduced by Morrey [10] in 1952.
@@
A real function $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{{n\times m}}\to\mathbb{R}$ is quasiconvex in the variable $\xi\in\mathbb{R}^{n\times m}$ if there exists a subset $Z$ of $\mathbb{R}^{n}$ with ${{\mathscr{L}}}^{n}(Z)=0$, such that for every ${x}\in\mathbb{R}^{n}\setminus Z$ and for every ${s}\in\mathbb{R}^{m}$ the function $\xi\mapsto f({x},{s},\xi)$ is quasiconvex.
More precisely, Morrey showed that under some strong regularity assumptions on the function $f$, the equivalence between its quasiconvexity and the sequential weak lower semicontinuity in $W^{1,\infty}(\Omega,\mathbb{R}^{m})$ of the functional
@@ 2. Quasiconvexity in the Riemannian setting
In this section, after introducing a suitable “Riemannian” definition of quasiconvexity generalizing the “classical” Euclidean one, we will prove that the sequential weak lower semicontinuity in $W^{1,\infty}$ of a functional holds if and only if the integrand is quasiconvex, according to such definition.
@@
(where the identification of $T_{x}M$ with ${{\mathbb{R}}}^{n}$ is done via an orthonormal basis of $T_{x}M$) and its diffeomorphic image
@@
We chose to adopt the notation $du[x]$ for the differential of a map $u:M\to{{\mathbb{R}}}^{m}$ at a point $x\in M$, for the sake of clarity in the computations that follow.
@@
Let $(M,g)$ be a smooth, connected and complete Riemannian manifold and $\mu$ its canonical volume measure. A continuous function $f:\mathscr{L}(TM,\mathbb{R}^{m})\to\mathbb{R}$ is called quasiconvex if for every $x_{0}\in M$, $\alpha_{x_{0}}\in(T_{x_{0}}M)^{m}\subseteq\mathscr{L}(TM,\mathbb{R}^{m})$ and $\varphi\in C^{\infty}_{c}(Q_{x_{0}}^{r},\mathbb{R}^{m})$, there holds
@@
is sequentially lower semicontinuous in the weak* topology of $W^{1,\infty}(\Omega,{{\mathbb{R}}}^{m})$, hence $\widetilde{f}$ is quasiconvex according to the Definition 1.1, by the Acerbi–Fusco Theorem 1.2.
@@ 3. Quasiconvexity and semicontinuity
We are going to show the analogues of some special cases of Theorem 1.2 of Acerbi and Fusco in our context, following the line of [2] (we mention that the main argument in the next Theorem 3.1 is due to Fonseca–Muller [5]). We underline (see the discussion above) that we are generalizing the special case of integrands which are continuous and depend only on $x$ and $Du$ (but not on $u$ – see the final section), from the Euclidean to the Riemannian setting.
@@
Let $n\in{{\mathbb{N}}}$ be the dimension of $M$. Let $x_{0}\in\Omega$ and we assume that $r>0$ is small enough for $Q_{x_{0}}^{r}$ to be contained in a neighbourhood $U\subseteq\Omega$ of $x_{0}$ such that the vector bundle $W^{1,\infty}(\Omega,{{\mathbb{R}}}^{m})$ can be “trivialized” as we discussed above at the beginning of this section, by choosing an orthonormal frame $E_{1},\dots,E_{n}$ in $U$, with an associated map $I$ and distance $\delta$ as in formulas (2.1) and (2.2). Moreover, since in what follows all the arguments of the continuous function $f$ will be bounded (since we are working in $W^{1,\infty}(\Omega,{{\mathbb{R}}}^{m})$ and $u_{j}$ is a bounded sequence in $W^{1,\infty}$, being weakly convergent), we can assume that $f$ is bounded and has a uniform modulus of continuity $\omega$ in $\mathscr{L}(TU,\mathbb{R}^{m})$ (with respect to the distance $\delta$) which is continuous, bounded and concave, hence subadditive. Moreover, since it is bounded, by possibly adding a constant to $f$, we can also assume that it is positive.
@@
where the function $o(1)$ is independent of $j\in{{\mathbb{N}}}$, while the functions $o_{j}(1)$ are equal to $\eta(r,\|d\varphi_{j}\|)$ for some function $\eta$ going to zero as $r\to 0$ and monotone nondecreasing (and we can also clearly assume continuous from the right) in its second argument.
@@
$\varphi_{h}\mathrel{\hbox{\set@color\hskip 4.93054pt\hskip-4.93054pt\hbox{\set@color$\xrightharpoonup{\ \ \,}$}\hskip-4.93054pt\hskip-3.71524pt\raisebox{-2.5pt}{\hbox{\set@color$\scriptstyle{\,\star\,}$}}\hskip-3.71524pt\hskip 4.93054pt}}0\quad\text{ in }W^{1,\infty}_{0}(Q_{x_{0}}^{r},\mathbb{R}^{m})$
@@
for every $r>0$ small enough. Then, arguing as in the proof of Theorem 3.1 (more precisely, following the same argument leading to estimate (3.13)), considering a suitable modulus of continuity $\omega$ of $f$, we have
@@ 4. Some remarks and possible research directions
A natural continuation of our work would be to extend the previous results also to the $W^{1,p}(\Omega,\mathbb{R}^{m})$ setting, with $p\in[1,+\infty)$, in order to obtain the analogue of Theorem 1.3 (notice that the fact that the semicontinuity of the functional implies quasiconvexity of $f$ follows immediately by Theorem 3.2 in such setting, as in the Euclidean case).
Focardi and Spadaro in [4] showed an extension of the Acerbi–Fusco theorems (with $f$ continuous, like us) in the case of Sobolev maps $u:\Omega\to M$, with $\Omega$ an open bounded subset of $\mathbb{R}^{n}$ and $M$ a Riemannian manifold (this case actually “corresponds” to the Euclidean situation with a continuous function $f$ depending on $u$ and $Du$, but not on $x$). Therefore, a possible future research line could be to “combine” Theorem 3.3 with the results of Focardi and Spadaro, thus obtaining a characterization of quasiconvexity in the completely Riemannian context of Sobolev maps between two Riemannian manifolds.
Moreover, the full generalization of the Acerbi–Fusco results would be to show them for the “naturally defined” Carathéodory functions on a Riemannian manifold, that is, functions which are Carathéodory in the usual way after expressing them by means of the local “trivializations” of the bundle $\mathscr{L}(TM,\mathbb{R}^{m})$ (or of $\mathscr{L}(TM,TN)$ after “combining” our definition with the one of Focardi–Spadaro), as we did at the beginning of this section.